OpenStudy (anonymous):

(3x-4)^3 Express as a single polynomial in standard form, and explain please!

OpenStudy (saifoo.khan):

Rewrite: (3x-4)(3x-4)(3x-4) Solve.

OpenStudy (anonymous):

just use this formula

OpenStudy (amistre64):

isnt (-b)^3 equal to - b^3 ?

OpenStudy (anonymous):

oh...sorry.... (a-b)^3 = a^3 -3a^2 b +3a b^2 - b^3

OpenStudy (amistre64):

:) better

OpenStudy (anonymous):

Can you explain how to write it out with the numbers in it?

OpenStudy (anonymous):

Should we bother mentioning Pascal's Triangle here?

OpenStudy (anonymous):

I think the best way to see this is to start with the expansion that saifoo showed and start by multiplying the first two binomials together and simplifying that, then multiply by the third binomial.

OpenStudy (anonymous):

27x^3 - 108x^2 +144x -64?

OpenStudy (anonymous):

Looks good!

OpenStudy (anonymous):

Do you see how that lines up with Yahoo's (a-b)^3 = a^3 -3a^2 b +3a b^2 - b^3 ?

OpenStudy (anonymous):

Actually, yeah! They're both useful

OpenStudy (anonymous):

Ok, now go research Pascal's Triangle and "binomial coefficients. You'll learn a lot of cool stuff.

OpenStudy (anonymous):

lol....that will make it ugly...use...that formula

OpenStudy (anonymous):

Heh heh, yes ugly, but in a beautiful way. He already simplified it correctly using good ol' distributive property, now it's time to 'expand' to bigger and better things. (pun definitely intended)